how to graph -4/3,1,-1.2,square root of 3,1.9

Answer by CubeyThePenguin(3113)

-4/3 = -1.333
1 = 1.000
-1.2 = -1.200
sqrt(3) = 1.732
1.9 = 1.900

-2..........-1..........0..........1.........2
<---|--|----------------|-----|--|----->